OpenStudy (anonymous):

Please help! The length of one leg of a right triangle is x - 6 centimeters, and the area is 1/2 x^2 - 7 + 24 square centimeters. What is the length of the other leg?

OpenStudy (dtan5457):

Diagram?

OpenStudy (dtan5457):

It really depends which legs you are describing..hypotenuse..etc

OpenStudy (dtan5457):

Actually..wait I don't think you need a diagram

OpenStudy (dtan5457):

The formula for area is( base time height)/2

OpenStudy (dtan5457):

x-6 times what /2= 1/2 x^2 - 7 + 24

OpenStudy (anonymous):

i have no idea

OpenStudy (dtan5457):

Multiply the area back by 2 to get x^2-14x+48 that factors into what?

OpenStudy (dtan5457):

(x-6) and..?

OpenStudy (anonymous):

is it (x-8) ?

OpenStudy (dtan5457):

Yup, the missing leg is (x-8) because (x-6)(x-8)=x^2-14x+48 divide that by 2 and you get the correct area. 1/2x^2-7x+24

OpenStudy (anonymous):

oooh okay :) thanks